OpenStudy (anonymous):

Please help! Event A: At least one die shows a 4 Event B: The sum of the dice is greater than 5 What is P(A)? What is P(B)? What is P(A and B)? What is P(A or B)? What is P(A|B)?

OpenStudy (goformit100):

Are you able to make the Solution set of 'Event A' and 'Event B' ?

OpenStudy (anonymous):

You mean a table of the sum of the dice?

OpenStudy (anonymous):

I might have gotten this wrong but I got P(A) = 11/36 P(B) = 13/18 P(A and B) = 1/4 P(A or B) = 7/9 P(A|B) = 9/26

OpenStudy (kropot72):

The sample space has 36 possible combinations of numbers. These can be set out in column form as follows: 6,6 5,6 4,6 3,6 2,6 1,6 6,5 5,5 4,5 3,5 2,5 1,5 6,4 5,4 4,4 3,4 2,4 1,4 6,3 5,3 4,3 3,3 2,3 1,3 6,2 5,2 4,2 3,2 2,2 1,2 6,1 5,1 4,1 3,1 2,1 1,1

OpenStudy (kropot72):

I confirm your results for P(A) and P(B).

OpenStudy (kropot72):

P(A and B) is the intersection of A and B: \[A \cap B=\frac{1}{4}\]

OpenStudy (kropot72):

P(A or B) is A union B: \[P(A \cup B)=P(A)+P(B)-P(A \cap B)=\frac{11}{36}+\frac{13}{18}-\frac{1}{4}=\frac{7}{9}\]

OpenStudy (kropot72):

I confirm your result for P(A|B)

OpenStudy (kropot72):

@notabakery Do you need any explanation?

OpenStudy (anonymous):

Nope, that makes a lot more sense! Thank you!

OpenStudy (kropot72):

You're welcome :)